### v1

Abstract
In this paper, we study second-order algorithms for solving nonconvex-strongly concave minimax problems, which have attracted much attention in recent years in many fields, especially in machine learning. We propose a gradient norm regularized trust region (GRTR) algorithm to solve nonconvex-strongly concave minimax problems, where the objective function of the trust region subproblem in each iteration uses a regularized version of the Hessian matrix, and the regularization coefficient and the radius of the ball constraint are proportional to the square root of the gradient norm. The iteration complexity of the proposed GRTR algorithm to obtain an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point is proved to be upper bounded by $\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-3/2})$, where $\rho$ and $\kappa$ are the Lipschitz constant of the Jacobian matrix and the condition number of the objective function respectively, which matches the best known iteration complexity of second-order methods for solving nonconvex-strongly concave minimax problems. We further propose a Levenberg-Marquardt algorithm with a gradient norm regularization coefficient and use the negative curvature direction to correct the iteration direction (LMNegCur), which does not need to solve the trust region subproblem at each iteration. We also prove that the LMNegCur algorithm achieves an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point within $\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-3/2})$ number of iterations. Numerical results show the efficiency of both proposed algorithms.
Keywords:
Trust region, Levenberg-Marquardt algorithm, nonconvex-strongly concave minimax problems, second-order algorithms
MSC:
90C47, 90C26, 90C30
1 Introduction
In this paper, we consider the following unconstrained minimax optimization problem:
$\min_{x\in\mathbb{R}^{n}}\max_{y\in\mathbb{R}^{m}}f(x,y),$ (1.1)
where $f(x,y):\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is a continuously differentiable function. Throughout this paper, we assume that the function $f(x,y)$ is $\mu$-strongly concave in $y$ for all $x\in\mathbb{R}^{n}$ but possibly nonconvex in $x$. Recently, the minimax optimization problem has stimulated the research interest of many international scholars and has broad applications in machine learning and data science, such as generative adversarial networks (GANs) Goodfellow2014GenerativeAN ; Arjovsky2017WassersteinGA , adversarial learning Sinha2018CertifiableDR , AUC maximization Hanley1982TheMA ; Ying2016StochasticOA , and robust optimization Ben-Tal2009RoubstOP ; Gao2016DistributionallyRS .
Three main types of optimization algorithms have been developed to solve the minimax problem (1.1), including zero-order, first-order, and second-order optimization algorithms, which exploit the function value, gradient, and Hessian information of the objective function, respectively. Compared with zero-order and first-order optimization algorithms, second-order optimization algorithms have received more attention due to their faster convergence speed. In addition, second-order optimization algorithms can handle saddle points and local minima more effectively, thereby increasing the probability of finding the global optimal solution. In this paper, we focus on the second-order optimization algorithms for solving (1.1).
In order to solve some convex-concave minimax optimization problems with special structures, various second-order algorithms have been proposed recently, such as implicit algorithms Monteiro2012IterationOA ; Ostroukhov2020TensorMF ; Bullins2022HigherMF ; Jiang2022GeneralizedOM and explicit algorithms Huang2022CubicRN ; Huang2022AnAR ; Jiang2024AdaptiveAO ; Lin2022ExplicitSM ; Nesterov2006bCubicRO . In Adil2022LineSM ; Lin2022PerseusAS , it is proved that the iteration complexity of the second-order algorithm for solving convex-concave minimax optimization problems is lower bounded by $\Omega(\epsilon^{-2/3})$. In addition, in order to eliminate the algorithm’s need for problem parameter information, such as the Lipschitz constant and the upper bound of the distance from the initial point to the optimal solution, some parameter-free second-order algorithms have been proposed in Jiang2024AdaptiveAO ; Liu2022RegularizedNM ; Wang2024ParameterFS .
However, there are relatively few studies on second-order algorithms for solving nonconvex minimax optimization problems (1.1). Existing second-order algorithms can be divided into two categories, i.e., cubic regularization type algorithms Luo2022FindingSS ; Chen2023ACR and trust region type algorithms Yao2024TwoTR . For the cubic regularization type algorithms, Luo et al. Luo2022FindingSS proposed a Minimax Cubic Newton (MCN) algorithm, which can find an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point with an iteration complexity of $\mathcal{O}(\epsilon^{-3/2})$. MCN performs accelerated gradient ascent method to update $y$, which is then used to estimate the gradient and Hessian of $P(x)=\max_{y\in\mathbb{R}^{m}}f(x,y)$ involved in the cubic regularization update for $x$. Chen et al. Chen2023ACR proposed a Cubic-LocalMinimax algorithm with the same iteration complexity as Luo2022FindingSS . For the trust region type algorithms, Yao and Xu Yao2024TwoTR first proposed a minimax trust region (MINIMAX-TR) algorithm, which can be viewed as an inexact trust region method with fixed trust region radius of $\mathcal{O}(\sqrt{\epsilon})$. Nevertheless, this strategy is relatively conservative, leading to poor numerical performance. To overcome this shortcoming, they further proposed a minimax trust region algorithm with contractions and expansions (MINIMAX-TRACE) algorithm. Both algorithms have the same iteration complexity result as that in Luo2022FindingSS ; Chen2023ACR . However, the MINIMAX-TRACE algorithm needs function information of $P(x)$ to calculate the ratio of actual-to-predicted reduction and additional assumption that the gradient of $P(x)$ is upper bounded by a constant.
1.1 Contributions
In this paper, we focus on developing a gradient norm regularized second-order algorithm to solve the nonconvex-strongly concave minimax optimization problems. We summarize our contributions as follows.
We first propose a gradient norm regularized trust region (GRTR) algorithm to solve the nonconvex-strongly concave minimax optimization problems, where the objective function of the trust region subproblem in each iteration adopts a regularized version of the Hessian matrix, and the regularization coefficient and the radius of the ball constraint are proportional to the root mean square of the gradient norm. It can be shown that the iteration complexity of the GRTR algorithm to obtain an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point is bounded by $\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-3/2})$, where $\tilde{\mathcal{O}}$ hides only absolute constants and logarithmic factors. Ignoring the logarithmic factor, this iteration complexity is consistent with the state-of-art existing second-order algorithms Luo2022FindingSS ; Chen2023ACR ; Yao2024TwoTR . In addition, compared with the MINIMAX-TR algorithm Yao2024TwoTR , our method does not need to always fix the trust region radius, but the trust region radius varies according to the gradient norm, which leads to better numerical performance. Compared with the MINIMAX-TRACE algorithm Yao2024TwoTR , we do not require additional function information and do not need to assume that the gradient has a uniform upper bound.
We further propose a Levenberg-Marquardt algorithm with a gradient norm regularization coefficient and use the negative curvature direction to correct the iteration direction (LMNegCur), which does not need to solve the trust region subproblem at each iteration. The LMNegCur algorithm can also achieve an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point within $\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-3/2})$ number of iterations.
1.2 Related Works
For solving convex-concave minimax optimization problems, there are currently a number of research results on first-order optimization algorithms, such as the gradient extrapolation algorithm Korpelevich1976TheEM , the mirror proximal algorithm Nemirovski2004ProxWR , the dual extrapolation algorithm Nesterov2007DualEA , and the optimistic gradient descent algorithm (OGDA) Popov1980AMO , which can achieve the optimal iteration complexity among algorithms that only use first-order information to solve convex-concave minimax problems, namely $\mathcal{O}(\epsilon^{-1})$Ouyang2021LowerCB .
There are also some existing research results on first-order algorithms for solving nonconvex-strongly concave minimax problems Bot ; Huang ; Jin ; Lin2019 ; Lin2020 ; Lu ; Rafique . To obtain an $\epsilon$-first-order stationary point of $P(x)$, or the stationary points of $f(x,y)$, these algorithms can achieve an iteration complexity of $\tilde{\mathcal{O}}\left(\kappa_{y}^{2}\epsilon^{-2}\right)$, where $\kappa_{y}$ is the condition number of $f(x,\cdot)$. In addition, Zhang et al. Zhang2021 can improve the iteration complexity to $\mathcal{O}\left(\sqrt{\kappa_{y}}\epsilon^{-2}\right)$ by introducing a general acceleration framework.
For general nonconvex-concave minimax problems, there are two types of algorithms, i.e., multi-loop algorithms Kong ; Lin2020 ; Nouiehed ; Ostro ; Rafique ; Thek2019 and single-loop algorithms Lin2019 ; Lu ; Xu ; Zhang ; Zheng2023 . The MINIMAX-PPA algorithm proposed in Lin2020 achieves the best known iteration complexity, i.e., $\tilde{\mathcal{O}}(\epsilon^{-2.5})$, among multi-loop algorithms for solving nonconvex-concave minimax problems. For solving nonconvex-concave minimax problems, the unified single-loop alternating gradient projection (AGP) algorithm Xu was the first single-loop algorithm to achieve an iteration complexity of $\mathcal{O}(\epsilon^{-4})$. Later, the smoothed GDA algorithm was also proposed Zhang , achieving the same iteration complexity. For solving the special nonconvex-linear minimax problems, Pan et al. Pan proposed a new alternating gradient projection algorithm with an iteration complexity of $\mathcal{O}(\epsilon^{-3})$. Very recently, Xu et al. XuJiangLiuSo proposed a riemannian alternating descent ascent algorithmic framework for nonconvex-linear minimax problems on riemannian manifolds, achieving the same complexity.
There are also some research results on the zeroth-order method for solving the black-box minimax problems, where we can only access function values. For the nonconvex-strongly concave setting, various algorithms have been proposed, eg., the ZO-Min-Max algorithm Liu , the zeroth-order gradient descent multi-step ascent algorithm Wang , and the accelerated zeroth-order momentum descent ascent algorithm Huang . Among these algorithms, Huang achieves the best known iteration complexity of $\tilde{\mathcal{O}}(\kappa_{y}^{4.5}\epsilon^{-3})$, where $\kappa_{y}$ is the condition number for $f(x,\cdot)$. For the nonconvex-concave setting, Xu et al. xu21zeroth proposed a zero-order alternating stochastic gradient projection (ZO-AGP) algorithm with aniteration complexity of $\mathcal{O}(\epsilon^{-4})$. For the nonconvex-nonconcave setting, Xu et al. xu22zeroth proposed a zeroth-order alternating gradient descent ascent (ZO-AGDA) algorithm and a zeroth-order variance reduced alternating gradient descent ascent (ZO-VRAGDA) algorithm under the deterministic and the stochastic setting, respectivelly. The iteration complexity of the ZO-AGDA algorithm and the ZO-VRAGDA algorithm have been proved to be $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-3})$, respectively.
[…]
2 A gradient norm regularized trust region algorithm
In this section, we propose a gradient norm regularized trust region (GRTR) algorithm for solving (1.1). The proposed algorithm is inspired by the universal trust-region (UTR) method Jiang2023AUT , which is a trust region variant for solving $\min_{x\in\mathbb{R}^{n}}P(x)$. In particular, at each iteration, the UTR method solves the following gradient-regularization trust-region subproblem:
$\begin{array}[]{l}\mathop{\min}\limits_{s}~{}\nabla P(x_{t})^{\top}s+\frac{1}{2}s^{\top}(\nabla^{2}P(x_{t})+\sigma_{t}\|\nabla P(x_{t})\|^{1/2}I)s\\ ~{}\text{s.t.}~{}~{}\|s\|\leqslant r_{t}\|\nabla P(x_{t})\|^{1/2},\end{array}$
where $\sigma_{t}$ and $r_{t}$ are iteration-dependent parameters, the objective function of the trust region subproblem uses a regularized version of the Hessian matrix, and the regularization coefficient and the radius of the ball constraint are proportional to the root mean square of the gradient norm. The proposed GRTR algorithm can be regarded as an inexact generalization version of UTR for solving (1.1) with $P(x):=\max_{y\in\mathbb{R}^{m}}f(x,y)$, which consists of the following two important algorithmic components at each iteration:
  • Gradient ascent update: for a given $x_{t}$, run $N_{t}$ iterations of gradient ascent step to obtain an approximated maximizer
    $y_{t}\approx y^{*}(x_{t})=\arg\max_{y\in\mathbb{R}^{m}}f(x,y).$
  • Trust region update: for given $(x_{t},y_{t})$, compute a direction $s_{t}$ such that it is a solution of the following trust-region subproblem:
    $\begin{array}[]{l}\mathop{\min}\limits_{s}~{}g_{t}^{\top}s+\frac{1}{2}s^{\top}(H_{t}+\sigma\|g_{t}\|^{1/2}I)s\\ ~{}\text{s.t.}~{}~{}\|s\|\leqslant r\max\{\|g_{t}\|^{1/2},\epsilon^{1/2}\},\end{array}$ (2.1)
    where $g_{t}=\nabla_{x}f(x_{t},y_{t})$ and $H_{t}=H(x_{t},y_{t})$ are inexact first-order and second-order information of $P(x)$ at $x_{t}$, respectively; $\sigma$ and $r$ are constants, which will be specified later.
The detailed algorithm is formally stated in Algorithm 1.
Algorithm 1 A gradient norm regularized trust region (GRTR) algorithm
  Step 1: Input $x_{0},y_{0},\eta_{y},\sigma,r,\epsilon,N_{t}$; Set $t=1$, $x_{1}=x_{0}$.
  Step 2: Update $y_{t}$:
   (a): Set $k=0$, $y_{k}^{t}=y_{t-1}$.
   (b): Update $y_{k}^{t}$:
$y_{k+1}^{t}=y_{k}^{t}+\eta_{y}\nabla_{y}f(x_{t},y_{k}^{t}).$ (2.2)
   (c): If $k\geqslant N_{t}-1$, set $y_{t}={y}_{N_{t}}^{t}$; otherwise, set $k=k+1,$ go to Step 2(b).
  Step 3: Set $g_{t}=\nabla_{x}f(x_{t},y_{t})$ and $H_{t}=H(x_{t},y_{t})$.
  Step 4: Update $s_{t}$ by solving the following trust region subproblem:
$\mathop{\min}\limits_{\|s\|\leqslant r\max\{\|g_{t}\|^{1/2},\epsilon^{1/2}\}}g_{t}^{\top}s+\frac{1}{2}s^{\top}(H_{t}+\sigma\|g_{t}\|^{1/2}I)s.$ (2.3)
  Step 5: Update $x_{t}$: $x_{t+1}=x_{t}+s_{t}$.
  Step 6: If some stationary condition or stopping criterion is satisfied, stop; otherwise, set $t=t+1,$ go to Step 2(a).
Note that in the trust region subproblem (2.3), the trust region radius varies with the gradient norm. When $\|g_{t}\|\leqslant\epsilon$, we fix the trust region radius to $\mathcal{O}(\sqrt{\epsilon})$. If we set $\sigma=0$ and fix the trust region radius to $\mathcal{O}(\sqrt{\epsilon})$, then the GRTR algorithm is equivalent to the MINIMAX-TR algorithm Yao2024TwoTR . Compared with the MINIMAX-TRACE algorithm Yao2024TwoTR , the GRTR algorithm does not need to calculate the function value of $P(x)$. In the next section, we will prove that the iteration complexity of the GRTR algorithm to solve (1.1) for an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$ is Note that in the trust region subproblem (2.3), the trust region radius varies with the gradient norm. When $\|g_{t}\|\leqslant\epsilon$, we fix the trust region radius to $\mathcal{O}(\sqrt{\epsilon})$. If we set $\sigma=0$ and fix the trust region radius to $\mathcal{O}(\sqrt{\epsilon})$, then the GRTR algorithm is equivalent to the MINIMAX-TR algorithm Yao2024TwoTR . Compared with the MINIMAX-TRACE algorithm Yao2024TwoTR , the GRTR algorithm does not need to calculate the function value of $P(x)$. In the next section, we will prove the iteration complexity of the GRTR algorithm to solve (1.1) for an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$ is $\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-1.5})$.
2.1 Complexity analysis
Before we prove the iteration complexity of the GRTR algorithm, we first make the following assumptions about $f(x,y)$ in this paper.
Assumption 2.1
$f(x,y)$ satifies the following assumptions:
  1. (i)
    $f(\cdot,\cdot)$ is $\ell$-Lipschitz continuous, i.e., for any $(x,y),(x^{\prime},y^{\prime})$, it holds
    $\|\nabla f(x,y)-\nabla f(x^{\prime},y^{\prime})\|\leqslant\ell\|(x,y)-(x^{\prime},y^{\prime})\|.$
  2. (ii)
    The Jacobian matrices $\nabla_{xx}^{2}f(x,y)$, $\nabla_{xy}^{2}f(x,y)$, $\nabla_{yx}^{2}f(x,y)$ and $\nabla_{yy}^{2}f(x,y)$ are $\rho$-Lipschitz continuous.
  3. (iii)
    $P(x)$ is bounded below by a constant $P^{*}$.
Next, similar to Luo2022FindingSS ; Chen2023ACR ; Yao2024TwoTR , the first-order and second-order stationary points are defined as follows.
Definition 2.1
Suppose that Assumption 2.1 holds. We call $x$ an $\epsilon$-first-order stationary point of $P(x)$ if $\|\nabla{P(x)}\|\leqslant\epsilon$.
Definition 2.2
Suppose that Assumption 2.1 holds. We call $x$ an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$ if $\|\nabla P(x)\|\leqslant\xi\epsilon$ and $\nabla^{2}P(x)\succeq-\theta\sqrt{\epsilon}I$, where $\xi$ and $\theta$ are positive constants independent of $\epsilon$.
Lemma 2.1 (Chen2023ACR )
Suppose that Assumption 2.1 holds. $P(x)$ has $L_{1}:=(\kappa+1)\ell$-Lipschitz continuous gradient with $\kappa=\ell/\mu$. Furthermore, $y^{*}(x)$ is well-defined and $\kappa$-Lipschitz continuous, and $\nabla P(x)=\nabla_{x}f(x,y^{*}(x))$.
Denote $H(x,y)$ $=$ $[\nabla_{xx}^{2}f-\nabla_{xy}^{2}f(\nabla_{yy}^{2}f)^{-1}\nabla_{yx}^{2}f](x,y)$.
Lemma 2.2 (Chen2023ACR )
Suppose that Assumption 2.1 holds. $H(x,y)$ is a Lipschitz continuous mapping with Lipschitz constant $L_{H}=\rho(1+\kappa)^{2}$. Additionally, the Hessian of $P(x)$ satisfies $\nabla^{2}P(x)=H(x,y^{*}(x))$, and it is Lipschitz continuous with constant $L_{2}=\rho(1+\kappa)^{3}$.
We first present some preliminary analysis of our method. The following lemma gives the optimality conditions for the trust region subproblem (2.3), which can be directly obtained from Theorem 4.1 in Jorge1999NumericalO and will not be repeated here.
Lemma 2.3
The direction $s_{t}$ is a solution of (2.3) if and only if there exists a dual multiplier $\lambda_{t}\geqslant 0$ such that
$\|s_{t}\|\leqslant r\max\{\|g_{t}\|^{1/2},\epsilon^{1/2}\},$ (2.4)
$\lambda_{t}(\|s_{t}\|-r\max\{\|g_{t}\|^{1/2},\epsilon^{1/2}\})=0,$ (2.5)
$(H_{t}+\sigma\|g_{t}\|^{1/2}I+\lambda_{t}I)s_{t}=-g_{t},$ (2.6)
$H_{t}+\sigma\|g_{t}\|^{1/2}I+\lambda_{t}I\succeq 0.$ (2.7)
In the remainder of this paper, we use $(s_{t},\lambda_{t})$ to denote the primal-dual solution pair of a subproblem at iteration $t$. The following lemma shows that by running a sufficient number of gradient ascent steps, we can obtain sufficiently accurate estimators of the gradient and Hessian.
Lemma 2.4 (Yao2024TwoTR )
Suppose that Assumption 2.1 holds. For any given $\epsilon_{1}>0$ and $\epsilon_{2}>0$, choose
$\eta_{y}=\frac{2}{\ell+\mu},~{}\begin{cases}N_{1}\geqslant\kappa\log\frac{\|y_{-1}-y^{*}(x_{0})\|}{A},\\ N_{t}\geqslant\kappa\log\frac{A+\kappa\|s_{t-1}\|}{A},\hskip 28.45274ptt\geqslant 2,\end{cases}$ (2.8)
where $A=\min\{\epsilon_{1}/l,\epsilon_{2}/L_{H}\}$, then $\|\nabla P(x_{t})-g_{t}\|\leqslant\epsilon_{1}$ and $\|\nabla^{2}P(x_{t})-H_{t}\|\leqslant\epsilon_{2}$.
From the optimality condition (2.4)-(2.7) and Lemma 2.4, we can prove the conclusions in the following lemma and give the key inequalities about the function and gradient in two different cases $\|g_{t}\|\geqslant\epsilon$ and $\|g_{t}\|\leqslant\epsilon$.
Lemma 2.5
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by Algorithm 1. Set
$\sigma=\frac{\sqrt{L_{2}}}{2},~{}r=\frac{1}{4\sqrt{L_{2}}},~{}\epsilon_{1}=\min\{\frac{1}{96},\frac{\sqrt{L_{2}}}{16L_{1}}\}\epsilon^{3/2},~{}\epsilon_{2}=\frac{\sqrt{L_{2}}}{12}\epsilon^{1/2},$ (2.9)
and choose $N_{t}$ in (2.8). When $\|g_{t}\|\geqslant\epsilon$, if $\lambda_{t}>0$, we have
$P(x_{t+1})\leqslant P(x_{t})-\frac{1}{128\sqrt{L_{2}}}\|g_{t}\|^{3/2},~{}\|g_{t+1}\|\leqslant 2+\frac{L_{1}}{4\sqrt{L_{2}}}\|g_{t}\|^{1/2}+\|g_{t}\|;\$ (2.10)
Otherwise if $\lambda_{t}=0$, we have
$P(x_{t+1})\leqslant P(x_{t}),~{}\|g_{t+1}\|\leqslant\frac{1}{3}\|g_{t}\|.$ (2.11)
Proof
By Lemma 2.2 and the Cauchy-Swartz inequality, we obtain
$~{}P(x_{t+1})-P(x_{t})$
$\leqslant$ $~{}\nabla P(x_{t})^{\top}s_{t}+\frac{1}{2}s_{t}^{\top}\nabla^{2}P(x_{t})s_{t}+\frac{L_{2}}{6}\|s_{t}\|^{3}$
$=$ $~{}g_{t}^{\top}s_{t}+\frac{1}{2}s_{t}^{\top}H_{t}s_{t}+(\nabla P(x_{t})-g_{t})^{\top}s_{t}+\frac{1}{2}s_{t}^{\top}(\nabla^{2}P(x_{t})-H_{t})s_{t}+\frac{L_{2}}{6}\|s_{t}\|^{3}$
$\leqslant$ $~{}g_{t}^{\top}s_{t}+\frac{1}{2}s_{t}^{\top}H_{t}s_{t}+\|\nabla P(x_{t})-g_{t}\|\|s_{t}\|+\frac{1}{2}\|\nabla^{2}P(x_{t})-H_{t}\|\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3}$
$\leqslant$ $~{}g_{t}^{\top}s_{t}+\frac{1}{2}s_{t}^{\top}H_{t}s_{t}+\epsilon_{1}\|s_{t}\|+\frac{\epsilon_{2}}{2}\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3},$ (2.12)
where the last inequality is by Lemma 2.4. By (2.6) and (2.7), it is easy to verify that
$g_{t}^{\top}s_{t}+\frac{1}{2}s_{t}^{\top}H_{t}s_{t}=$ $-\frac{1}{2}s_{t}^{\top}(H_{t}+\sigma\|g_{t}\|^{1/2}+\lambda_{t}I)s_{t}-\frac{1}{2}s_{t}^{\top}(\sigma\|g_{t}\|^{1/2}+\lambda_{t}I)s_{t}$
$\leqslant$ $-\frac{1}{2}s_{t}^{\top}(\sigma\|g_{t}\|^{1/2}+\lambda_{t}I)s_{t}.$ (2.13)
Combining (2.1) and (2.1), we get
$P(x_{t+1})\leqslant P(x_{t})-\frac{1}{2}s_{t}^{\top}(\sigma\|g_{t}\|^{1/2}+\lambda_{t}I)s_{t}+\epsilon_{1}\|s_{t}\|+\frac{\epsilon_{2}}{2}\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3}.$ (2.14)
For the case $\|g_{t}\|\geqslant\epsilon$ and $\lambda_{t}>0$, the complementary property (2.5) implies that
$\|s_{t}\|=r\|g_{t}\|^{1/2}.$ (2.15)
By plugging (2.15) and (2.9) into (2.14), we then obtain
$P(x_{t+1})-P(x_{t})\leqslant$ $-\frac{1}{2}\sigma r^{2}\|g_{t}\|^{3/2}+\frac{1}{96}r\epsilon\|g_{t}\|^{1/2}+\frac{\sqrt{L_{2}}}{24}r^{2}\epsilon^{1/2}\|g_{t}\|+\frac{L_{2}}{6}r^{3}\|g_{t}\|^{3/2}$
$\leqslant$ $-\frac{1}{128\sqrt{L_{2}}}\|g_{t}\|^{3/2},$ (2.16)
which shows the first inequaity of (2.10). By the Cauchy-Swartz inequality, Lemmas 2.1 and 2.4, we get
$\|g_{t+1}\|\leqslant$ $~{}\|g_{t+1}-\nabla P(x_{t+1})\|+\|\nabla P(x_{t+1})-\nabla P(x_{t})\|+\|\nabla P(x_{t})-g_{t}\|+\|g_{t}\|$
$\leqslant$ $~{}2\epsilon_{1}+L_{1}\|s_{t}\|+\|g_{t}\|.$ (2.17)
Plugging (2.15) and (2.9) into (2.1), the second inequaity of (2.10) is proved.
Next, we consider the case $\|g_{t}\|\geqslant\epsilon$ and $\lambda_{t}=0$. Combining (2.14) with $\|s_{t}\|\leqslant r\|g_{t}\|^{1/2}$ and (2.9), we have
$~{}P(x_{t+1})-P(x_{t})$
$\leqslant$ $-(\frac{\sigma}{2}-\frac{L_{2}r}{6})\|g_{t}\|^{1/2}\|s_{t}\|^{2}+\min\{\frac{1}{96},\frac{\sqrt{L_{2}}}{16L_{1}}\}\epsilon^{3/2}\|s_{t}\|+\frac{\sqrt{L_{2}}}{24}\epsilon^{1/2}\|s_{t}\|^{2}$
$\leqslant$ $-(\frac{\sqrt{L_{2}}}{6}-\min\{\frac{1}{96},\frac{\sqrt{L_{2}}}{16L_{1}}\}\frac{\epsilon^{3/2}}{\|g_{t}\|^{1/2}\|s_{t}\|})\|g_{t}\|^{1/2}\|s_{t}\|^{2}.$ (2.18)
By (2.6) and $\|H_{t}\|\leqslant\ell(1+\kappa)=L_{1}$, we can easily get
$\|g_{t}\|\leqslant(L_{1}+\sigma\|g_{t}\|^{1/2})\|s_{t}\|,$
which implies that
$~{}\epsilon^{2/3}\leqslant\epsilon^{1/2}\|g_{t}\|\leqslant(L_{1}+\sigma)\|g_{t}\|^{1/2}\|s_{t}\|.$ (2.19)
Then, the first inequaity of (2.11) holds by combining (2.1) with (2.19). By the Cauchy-Swartz inequality, Lemmas 2.2 and 2.4, (2.6) and $\|s_{t}\|\leqslant r\|g_{t}\|^{1/2}$, we have
$\|g_{t+1}\|\leqslant$ $~{}\|\nabla P(x_{t+1})-g_{t+1}\|+\|\nabla P(x_{t+1})-\nabla P(x_{t})-\nabla^{2}P(x_{t})s_{t}\|$
$~{}+\|\nabla P(x_{t})-g_{t}\|+\|\nabla^{2}P(x_{t})-H_{t}\|\|s_{t}\|+\|g_{t}+H_{t}s_{t}\|$
$\leqslant$ $~{}2\epsilon_{1}+\frac{L_{2}}{2}\|s_{t}\|^{2}+\epsilon_{2}\|s_{t}\|+\sigma\|g_{t}\|^{1/2}\|s_{t}\|$
$\leqslant$ $~{}(2\frac{\epsilon_{1}}{\|g_{t}\|}+\frac{L_{2}}{2}r^{2}+\frac{\epsilon_{2}}{\|g_{t}\|^{1/2}}r+\sigma r)\|g_{t}\|.$ (2.20)
The second inequaity of (2.11) follows by plugging (2.9) into (2.1).
Lemma 2.6
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by Algorithm 1 with parameter settings in (2.9). Then for the case $\|g_{t}\|\leqslant\epsilon$, if $\lambda_{t}\leqslant\sqrt{L_{2}\epsilon}$, we have already found an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point $x_{t}$; if $\lambda_{t}\geqslant\sqrt{L_{2}\epsilon}$, we have
$P(x_{t+1})-P(x_{t})\leqslant-\frac{1}{128}\sqrt{L_{2}}\epsilon^{3/2},~{}\|g_{t+1}\|\leqslant 2+\frac{L_{1}}{4\sqrt{L_{2}}}.$ (2.21)
Proof
We first consider the case $\|g_{t}\|\leqslant\epsilon$ and $\lambda_{t}\leqslant\sqrt{L_{2}\epsilon}$. By $\|g_{t}\|\leqslant\epsilon$, Lemma 2.4 and (2.9), we have
$\|\nabla P(x_{t})\|\leqslant\|\nabla P(x_{t})-g_{t}\|+\|g_{t}\|\leqslant\epsilon_{1}+\|g_{t}\|\leqslant\frac{97}{96}\epsilon.$ (2.22)
By (2.7), Lemma 2.4 and (2.9), we have
$\nabla^{2}P(x_{t})\succeq H_{t}-\|\nabla^{2}P(x_{t})-H_{t}\|I\succeq-\lambda_{t}I-\sigma\|g_{t}\|^{1/2}I-\epsilon_{2}I\succeq-\frac{19}{12}\sqrt{L_{2}\epsilon}.$ (2.23)
Then, (2.22) and (2.23) shows that $x_{t}$ is an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point.
For the case $\|g_{t}\|\leqslant\epsilon$ and $\lambda_{t}\geqslant\sqrt{L_{2}\epsilon}$, by (2.14), it follows that
$P(x_{t+1})-P(x_{t})\leqslant$ $-\frac{1}{2}\lambda_{t}\|s_{t}\|^{2}+\epsilon_{1}\|s_{t}\|+\frac{\epsilon_{2}}{2}\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3}$
$\leqslant$ $-\frac{1}{4}\sqrt{L_{2}\epsilon}\|s_{t}\|^{2}+\epsilon_{1}\|s_{t}\|+\frac{\epsilon_{2}}{2}\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3}.$
The rest of the proof is almost the same as in (2.1) and (2.1) by replacing $\|g_{t}\|$ with $\epsilon$. We omit the details here for simplicity.
We are now ready to establish the iteration complexity for the GRTR algorithm. In particular, let $\epsilon>0$ be any given target accuracy. We denote the first iteration index to achieve an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point by
${T}(\epsilon):=\min\{t|\|\nabla P(x_{t})\|\leqslant\xi\epsilon~{}\text{and}~{}\nabla^{2}P(x_{t})\succeq-\theta\sqrt{\epsilon}I\},$ (2.24)
where $\xi$ and $\theta$ are some constants independent of $\epsilon$. We further denote the first iteration index of the GRTR algorithm to achieve $\|g_{t}\|\leqslant\epsilon$ and $\lambda_{t}\leqslant\sqrt{L_{2}\epsilon}$ by
$\tilde{T}(\epsilon):=\min\{t|\|g_{t}\|\leqslant\epsilon~{}\text{and}~{}\lambda_{t}\leqslant\sqrt{L_{2}\epsilon}\}.$
From Lemma 2.6, it can be known that $x_{\tilde{T}(\epsilon)}$ is an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point, indicating that $T(\epsilon)\leqslant\tilde{T}(\epsilon)$.
Let us define the following index sets to facilitate the complexity analysis:
$\mathcal{F}_{\tilde{T}(\epsilon)}=\big{\{}t<\tilde{T}(\epsilon):P(x_{t+1})-P(x_{t})\leqslant-\frac{1}{128\sqrt{L_{2}}}\max\{\|g_{t}\|^{3/2},\epsilon^{3/2}\}\big{\}},$ (2.25)
$\mathcal{G}_{\tilde{T}(\epsilon)}=\left\{t<\tilde{T}(\epsilon):\|g_{t+1}\|\leqslant\frac{1}{3}\|g_{t}\|,t\notin\mathcal{F}_{\tilde{T}(\epsilon)}\right\}.$ (2.26)
According to Lemmas 2.5 and 2.6, each iteration $t$ belongs to one of the two sets $\mathcal{F}_{\tilde{T}(\epsilon)}$ and $\mathcal{G}_{\tilde{T}(\epsilon)}$. Therefore, our aim is to establish an upper bound for the cardinality of sets $\mathcal{F}_{\tilde{T}(\epsilon)}$ and $\mathcal{G}_{\tilde{T}(\epsilon)}$.
In the following lemma, we analyze $\mathcal{F}_{\tilde{T}(\epsilon)}$ by evaluating the decrease in function value.
Lemma 2.7
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by Algorithm 1 with parameter settings in (2.9). We have
$|\mathcal{F}_{\tilde{T}(\epsilon)}|\leqslant 128\sqrt{L_{2}}(P(x_{0})-P^{*})\epsilon^{-3/2}.$ (2.27)
Proof
According to Lemmas 2.5 and 2.6, for any $t\geqslant 1$, we have
$P(x_{t+1})\leqslant P(x_{t}).$ (2.28)
In addition, for any $t\in\mathcal{F}_{\tilde{T}(\epsilon)}$, we can easily obtain that
$\frac{1}{128\sqrt{L_{2}}}\epsilon^{3/2}\leqslant\frac{1}{128\sqrt{L_{2}}}\max\{\|g_{t}\|^{3/2},\epsilon^{3/2}\}\leqslant P(x_{t})-P(x_{t+1}).$ (2.29)
From (2.28), (2.29) and Assumption 2.1, it follows that
$|\mathcal{F}_{\tilde{T}(\epsilon)}|\frac{1}{128\sqrt{L_{2}}}\epsilon^{3/2}\leqslant$ $\sum_{t\in\mathcal{F}_{\tilde{T}(\epsilon)}}(P(x_{t})-P(x_{t+1}))$
$\leqslant$ $\sum_{t=1}^{\tilde{T}(\epsilon)}(P(x_{t})-P(x_{t+1}))\leqslant P(x_{0})-P^{*},$
which completes the proof.
Now, it remains to establish an upper bound on the index set $|\mathcal{G}_{\tilde{T}(\epsilon)}|$.
Lemma 2.8
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by Algorithm 1 with parameter settings in (2.9). We have
$|\mathcal{G}_{\tilde{T}(\epsilon)}|\leqslant\lceil\log(G/\epsilon)/\log(3)\rceil|\mathcal{F}_{\tilde{T}(\epsilon)}|,$ (2.30)
where $G=\max\big{\{}2+\frac{L_{1}}{4\sqrt{L_{2}}}C^{1/2}+C,2+\frac{L_{1}}{4\sqrt{L_{2}}}\big{\}}$ with $C=\left(128\sqrt{L_{2}}\left(P(x_{0})-P^{*}\right)\right)^{2/3}=\mathcal{O}(\rho^{1/3}\kappa)$, which further implies $G=\mathcal{O}(\rho^{1/3}\kappa+\ell\rho^{-1/3})$.
Proof
We first establish the uniform upper bound of the norm of gradient. For any $t\in\mathcal{F}_{\tilde{T}(\epsilon)}$, we can conclude that
$\|g_{t}\|\leqslant C=\left(128\sqrt{L_{2}}\left(P(x_{0})-P^{*}\right)\right)^{2/3}$ (2.31)
by contradiction. Suppose that (2.31) does not hold. Then by $t\in\mathcal{F}_{\tilde{T}(\epsilon)}$, (2.28) and (2.29), we have
$P(x_{t+1})\leqslant P(x_{t})-\frac{1}{128\sqrt{L_{2}}}\|g_{t}\|^{3/2}\leqslant P(x_{0})-\frac{1}{128\sqrt{L_{2}}}\|g_{t}\|^{3/2}<P^{*},$
which is a contradiction. Combining (2.31) with (2.10) and (2.21), for any $t\in\mathcal{F}_{\tilde{T}(\epsilon)}$, we obtain that
$\|g_{t+1}\|\leqslant G=\max\big{\{}2+\frac{L_{1}}{4\sqrt{L_{2}}}C^{1/2}+C,2+\frac{L_{1}}{4\sqrt{L_{2}}}\big{\}}.$
Note that for any $t\in\mathcal{G}_{\tilde{T}(\epsilon)}$, the norm of the gradient will decrease linearly with factor $\frac{1}{3}$. We then deduce that the norm of gradient is uniformly bounded by $G$.
Next we give an upper bound on the index set $|\mathcal{G}_{\tilde{T}(\epsilon)}|$. Denote the maximum number of consecutive iterates in set $\mathcal{G}_{\tilde{T}(\epsilon)}$ as $n$. We prove $n\leqslant\lceil\log(G/\epsilon)/\log(3)\rceil$ by contradiction. Assume $n>\log(G/\epsilon)/\log(3)$. From $t\in\mathcal{G}_{\tilde{T}(\epsilon)}$ and $\|g_{t}\|\leqslant G$, we know that
$\|g_{t+n}\|\leqslant(\frac{1}{3})^{n}\|g_{t}\|\leqslant(\frac{1}{3})^{n}G<\epsilon,$
Then according to Lemma 2.6, $t+n\in\mathcal{F}_{\tilde{T}(\epsilon)}$, which contradicts the definition of $n$. Therefore, $n\leqslant\lceil\log(G/\epsilon)/\log(3)\rceil$ and the inequality (2.30) follows directly.
Next, we can give results for the iteration complexity.
Theorem 2.1
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be the sequence generated by Algorithm 1 with all parameters chosen the same as in Lemma 2.5. We have
$T(\epsilon)\leqslant 128(\lceil\log(G/\epsilon)/\log(3)\rceil+1)\sqrt{L_{2}}(P(x_{0})-P^{*})\epsilon^{-3/2}+1,$ (2.32)
where $G$ is defined as in Lemma 2.8.
The conclusion of the theorem can be directly proved by Lemmas 2.7 and 2.8, and the details are not elaborated here.
Remark 1
Theorem 2.1 shows that the iteration complexity of Algorithm 1 for finding a $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$ is $\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-1.5})$. This is consistent with the best iteration complexity of the second-order algorithm Luo2022FindingSS ; Chen2023ACR ; Yao2024TwoTR for solving non-convex-strongly concave minimax problems, omitting the logarithmic factor, and is better than the best iteration complexity of the first-order algorithm. In Yang2023AccI , a first-order algorithm is proposed to obtain the second-order stationary point with an iteration complexity of $\tilde{\mathcal{O}}(\kappa^{1.75}\epsilon^{-1.75})$. Compared with this algorithm, Algorithm 1 converges slightly faster, but each step requires solving the gradient regularization trust region subproblem.
Moreover, we can have a rough estimate of the number of inner iterations, i.e., $N_{t}$. If $x_{t}$ is not an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point, according to Lemmas 2.5 and 2.6, we conclude that either $\|s_{t}\|=r\max\{\|g_{t}\|^{1/2},\epsilon^{1/2}\}$ or $\|g_{t}\|\leqslant(L_{1}+\sigma\|g_{t}\|^{1/2})\|s_{t}\|$ for the case $\lambda_{t}=0$ and $\|g_{t}\|\geqslant\epsilon$. Then it follows that
$\|s_{t}\|\geqslant\mathcal{O}(\min\{\sqrt{\frac{\epsilon}{L_{2}}},\frac{\epsilon}{L_{1}+\sqrt{L_{2}}}\}).$
By further combining with Lemma 2.4, we obtain that $A=\mathcal{O}(\rho^{0.5}\kappa^{0.5}\ell^{-2}\epsilon^{1.5}+\rho^{-0.5}\kappa^{-0.5}\epsilon)$, $N_{1}=\mathcal{O}(\kappa\log\frac{\|y_{0}-y^{*}(x_{0})\|}{A})$ and $N_{t}=\mathcal{O}(\kappa\log\frac{A+\kappa\|s_{t-1}\|}{A})$ ($t\geqslant 2$). If the Nesterov’s accelerated gradient algorithm is used to solve the subproblem regarding $y$ as in Luo2022FindingSS , then the number of inner iterations will reduce to $N_{1}=\mathcal{O}(\sqrt{\kappa}\log\frac{\sqrt{\kappa+1}\|y_{0}-y^{*}(x_{0})\|}{A})$ and $N_{t}=\mathcal{O}(\sqrt{\kappa}\log(\sqrt{\kappa+1}+\frac{\sqrt{\kappa+1}\kappa\|s_{t-1}\|}{A})$ ($t\geqslant 2$) by Theorem 2 in Luo2022FindingSS .
3 A Levenberg-Marquardt algorithm with negative curvature correction
In the classic nonlinear minimization optimization, when the Hessian matrix is not positive definite, the negative curvature direction can be used to correct the iteration direction of the algorithm to converge to a second-order stationary point Gratton2023YetAF . In this section, we further propose a Levenberg-Marquardt (LM) algorithm with a gradient norm regularization coefficient Mischenko2023RegularizedNM , which uses the negative curvature direction to correct the iteration direction (LMNegCur) to solve (1.1) without solving the trust region subproblem at each iteration. The LMNegCur algorithm consists of the following two important algorithmic components at each iteration:
  • Gradient ascent update: for given $x_{t}$, run $N_{t}$ iterations of gradient ascent step to obtain an approximated maximizer
    $y_{t}\approx y^{*}(x_{t})=\arg\max_{y\in\mathbb{R}^{m}}f(x,y).$
  • LM or negative curvature update: for a given pair $(x_{t},y_{t})$, if $\lambda_{\min}(H_{t})\leqslant-\frac{\sqrt{L_{2}\max\{\|g_{t}\|,\epsilon\}}}{2}$, perform a negative curvature update
    $s_{t}=\sqrt{\frac{\max\{\|g_{t}\|,\epsilon\}}{L_{2}}}u_{t},~{}\text{where}~{}g_{t}^{\top}u_{t}\leqslant 0,\|u_{t}\|=1,H_{t}u_{t}=\lambda_{\min}(H_{t})u_{t};$
    otherwise, if $\|g_{t}\|\geqslant\epsilon$ and $\lambda_{\min}(H_{t})\geqslant-\dfrac{1}{2}\sqrt{L_{2}\|g_{t}\|}$, perform LM update with a gradient norm regularization coefficient
    $s_{t}=-(H_{t}+\sqrt{L_{2}\|g_{t}\|}I)^{-1}g_{t},$
    where $g_{t}=\nabla_{x}f(x_{t},y_{t})$ and $H_{t}=H(x_{t},y_{t})$ are inexact first-order and second-order information of $P(x)$ at $x_{t}$, respectively.
The detailed algorithm is formally stated in Algorithm 2.
Note that for the case $\|g_{t}\|\leqslant\epsilon,~{}\lambda_{\min}(H_{t})\geqslant-\frac{1}{2}\sqrt{L_{2}\epsilon}$, we have already achieved an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$, which will be proved later. The smallest eigenvalue of $H_{t}$ and the corresponding curvature direction can be retrieved via a Lanczos procedure. Since $\|H_{t}\|\leqslant L_{1}$, it was shown in Royer2018ComplexityAO (Lemma 9) that after at most $n$ (the dimension of $x$) iterations, the procedure obtains a unit vector $u$ such that $u^{\top}H_{t}u=\lambda_{\min}(H_{t})$ with probability 1. Therefore, using the smallest eigenvalue of the Hessian does not incur much cost.
Algorithm 2 A LMNegCur Algorithm
  Step 1: Input $x_{0},y_{0},\eta_{y},N_{t},\epsilon$; Set $t=1$, $x_{1}=x_{0}$.
  Step 2: Update $y_{t}$:
   (a): Set $k=0$, $y_{k}^{t}=y_{t-1}$.
   (b): Update $y_{k}^{t}$:
$y_{k+1}^{t}=y_{k}^{t}+\eta_{y}\nabla_{y}f(x_{t},y_{k}^{t}).$ (3.1)
   (c): If $k\geqslant N_{t}-1$, set $y_{t}={y}_{N_{t}}^{t}$; otherwise, set $k=k+1,$ go to Step 2(b).
  Step 3: If $\lambda_{\min}(H_{t})\leqslant-\frac{1}{2}\sqrt{L_{2}\max\{\|g_{t}\|,\epsilon\}}$,
$s_{t}=\sqrt{\frac{\max\{\|g_{t}\|,\epsilon\}}{L_{2}}}u_{t},~{}\text{where}~{}g_{t}^{\top}u_{t}\leqslant 0,\|u_{t}\|=1,H_{t}u_{t}=\lambda_{\min}(H_{t})u_{t};$ (3.2)
      otherwise, go to Step 4.
  Step 4: If $\|g_{t}\|\geqslant\epsilon$,
$s_{t}=-(H_{t}+\sqrt{L_{2}\|g_{t}\|}I)^{-1}g_{t};$ (3.3)
      otherwise, stop and output $(x_{t},y_{t})$.
  Step 5: Update $x_{t}$: $x_{t+1}=x_{t}+s_{t}.$ Set $t=t+1,$ go to Step 2(a).
In the following subsection, we will establish the iteration complexity of the LMNegCur algorithm to obtain an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$.
3.1 Complexity analysis
In the following lemmas, we first prove some key inequalities about functions and gradients in different cases, namely, Case 1 ($\|g_{t}\|\geqslant\epsilon$ and $\lambda_{\min}(H_{t})\geqslant-\dfrac{1}{2}\sqrt{L_{2}\|g_{t}\|}$), Case 2 ($\lambda_{\min}(H_{t})\leqslant-\dfrac{1}{2}\sqrt{L_{2}\max\{\|g_{t}\|,\epsilon\}}$), and Case 3 ($\|g_{t}\|\leqslant\epsilon$ and $\lambda_{\min}(H_{t})\geqslant-\dfrac{1}{2}\sqrt{L_{2}\epsilon}$).
Lemma 3.1
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by Algorithm 2. By choosing
$\epsilon_{1}=\min\{\frac{1}{36},\frac{\sqrt{L_{2}}}{12L_{1}}\}\epsilon^{3/2},~{}\epsilon_{2}=\frac{\sqrt{L_{2}}}{18}\epsilon^{1/2},$ (3.4)
in Lemma 2.4 and choosing $N_{t}$ in (2.8). For the case $\|g_{t}\|\geqslant\epsilon$, if $\lambda_{\min}(H_{t})\geqslant-\dfrac{1}{2}\sqrt{L_{2}\|g_{t}\|}$, we have
$~{}\|g_{t+1}\|$ $\leqslant\frac{7}{3}\|g_{t}\|,$ (3.5)
$P(x_{t+1})$ $\leqslant P(x_{t})-\dfrac{5}{18}\sqrt{L_{2}\|g_{t}\|}\bigg{(}\frac{-\sqrt{L_{2}\|g_{t}\|}+\sqrt{L_{2}\|g_{t}\|+72L_{2}\|g_{t+1}\|}}{6L_{2}}\bigg{)}^{2};$ (3.6)
If $\|g_{t+1}\|\geqslant\frac{1}{2}\|g_{t}\|$ holds, then we further have
$P(x_{t+1})\leqslant P(x_{t})-\frac{7}{36\sqrt{L_{2}}}\|g_{t}\|^{3/2}.$ (3.7)
Proof
By the same proof as in (2.1), we have
$P(x_{t+1})-P(x_{t})\leqslant g_{t}^{\top}s_{t}+\frac{1}{2}s_{t}^{\top}H_{t}s_{t}+\epsilon_{1}\|s_{t}\|+\frac{\epsilon_{2}}{2}\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3}.$
By plugging (3.3) into the above relation and using the assumption that $\lambda_{\min}(H_{t})\geqslant-\dfrac{1}{2}\sqrt{L_{2}\|g_{t}\|}$, we obtain that
$~{}P(x_{t+1})-P(x_{t})$
$\leqslant$ $-\frac{1}{2}s_{t}^{\top}(H_{t}+\frac{1}{2}\sqrt{L_{2}\|g_{t}\|})s_{t}-\frac{3}{4}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|^{2}+\epsilon_{1}\|s_{t}\|+\frac{\epsilon_{2}}{2}\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3}$
$\leqslant$ $-\frac{3}{4}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|^{2}+\epsilon_{1}\|s_{t}\|+\frac{\epsilon_{2}}{2}\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3}.$ (3.8)
To further bound the right-hand side of (3.1), we first give some bounds on $\|s_{t}\|$. By combining (3.3) with $\lambda_{\min}(H_{t})\geqslant-\dfrac{1}{2}\sqrt{L_{2}\|g_{t}\|}$ and $\|H_{t}\|\leqslant\ell(1+\kappa)=L_{1}$, we then obtain that
$\frac{\|g_{t}\|}{L_{1}+\sqrt{L_{2}\|g_{t}\|}}\leqslant\|s_{t}\|\leqslant\|(H_{t}+\sqrt{L_{2}\|g_{t}\|}I)^{-1}\|\|g_{t}\|\leqslant 2\sqrt{\dfrac{\|g_{t}\|}{L_{2}}},$ (3.9)
which implies that
$(L_{1}+\sqrt{L_{2}})\sqrt{\|g_{t}\|}\|s_{t}\|\geqslant\epsilon^{1/2}\|g_{t}\|\geqslant\epsilon^{3/2}.$ (3.10)
Combining (3.1) with (3.9) and (3.10), and by (3.4), we conclude that
$~{}P(x_{t+1})-P(x_{t})$
$\leqslant$ $~{}-\frac{5}{12}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|^{2}+\min\{\frac{1}{36},\frac{\sqrt{L_{2}}}{12L_{1}}\}\epsilon^{3/2}\|s_{t}\|+\frac{\sqrt{L_{2}}}{36}\epsilon^{1/2}\|s_{t}\|^{2}$
$\leqslant$ $-(\frac{7\sqrt{L_{2}}}{18}-\min\{\frac{1}{36},\frac{\sqrt{L_{2}}}{12L_{1}}\}\frac{\epsilon^{3/2}}{\sqrt{\|g_{t}\|}\|s_{t}\|})\sqrt{\|g_{t}\|}\|s_{t}\|^{2},$
$\leqslant$ $~{}-\dfrac{5}{18}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|^{2}.$ (3.11)
By (3.3) and using an argument similar to the proof of (2.1), we have
$\|g_{t+1}\|\leqslant 2\epsilon_{1}+\frac{L_{2}}{2}\|s_{t}\|^{2}+\epsilon_{2}\|s_{t}\|+\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|.$
By further combining the above relation with (3.4) and (3.10), we obtain that
$\|g_{t+1}\|\leqslant\frac{1}{6}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|+\frac{L_{2}}{2}\|s_{t}\|^{2}.$ (3.12)
Then, (3.5) holds directly by combining (3.12) with (3.9). On the other hand, (3.12) implies that
$\|s_{t}\|\geqslant\frac{-\sqrt{L_{2}\|g_{t}\|}+\sqrt{L_{2}\|g_{t}\|+72L_{2}\|g_{t+1}\|}}{6L_{2}}.$ (3.13)
Then, (3.6) holds directly by combining (3.1) with (3.13), and (3.7) holds by further combining (3.6) with $\|g_{t+1}\|\geqslant\frac{1}{2}\|g_{t}\|$. The proof is completed.
Lemma 3.2
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by Algorithm 2 with parameters chosen as in (3.4). For the case $\lambda_{\min}(H_{t})\leqslant-\dfrac{1}{2}\sqrt{L_{2}\max\{\|g_{t}\|,\epsilon\}}$, we have
$P(x_{t+1})\leqslant P(x_{t})-\frac{1}{36\sqrt{L_{2}}}\max\{\|g_{t}\|,\epsilon\}^{3/2},$ (3.14)
$\|g_{t+1}\|\leqslant 1+\frac{L_{1}}{\sqrt{L_{2}}}\sqrt{\max\{\|g_{t}\|,\epsilon\}}+\max\{\|g_{t}\|,\epsilon\}.$ (3.15)
Proof
By (3.2) and $\lambda_{\min}(H_{t})\leqslant-\dfrac{1}{2}\sqrt{L_{2}\max\{\|g_{t}\|,\epsilon\}}$, we have
$\|s_{t}\|=\sqrt{\frac{\max\{\|g_{t}\|,\epsilon\}}{L_{2}}},~{}g_{t}^{\top}s_{t}\leqslant 0,~{}g_{t}^{\top}s_{t}+\frac{1}{2}s_{t}^{\top}H_{t}s_{t}\leqslant-\frac{1}{4\sqrt{L_{2}}}\max\{\|g_{t}\|,\epsilon\}^{3/2}.$ (3.16)
Combining (2.1) and (2.1) with (3.16), and by (3.4), we obtain
$P(x_{t+1})-P(x_{t})\leqslant$ $-\frac{1}{12\sqrt{L_{2}}}\max\{\|g_{t}\|,\epsilon\}^{3/2}+\epsilon_{1}\sqrt{\frac{\max\{\|g_{t}\|,\epsilon\}}{L_{2}}}+\frac{\epsilon_{2}}{2}\frac{\max\{\|g_{t}\|,\epsilon\}}{L_{2}}$
$\leqslant$ $-\frac{1}{36\sqrt{L_{2}}}\max\{\|g_{t}\|,\epsilon\}^{3/2},$
and
$\|g_{t+1}\|\leqslant$ $~{}2\epsilon_{1}+L_{1}\|s_{t}\|+\|g_{t}\|\leqslant 1+\frac{L_{1}}{\sqrt{L_{2}}}\sqrt{\max\{\|g_{t}\|,\epsilon\}}+\max\{\|g_{t}\|,\epsilon\}.$
The proof is then completed.
Lemma 3.3
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by Algorithm 2 with parameters choosen as in (3.4). For the case $\|g_{t}\|\leqslant\epsilon$ and $\lambda_{\min}(H_{t})\geqslant-\dfrac{1}{2}\sqrt{L_{2}\epsilon}$, we have already found an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point $x_{t}$.
Proof
By $\|g_{t}\|\leqslant\epsilon$, $\lambda_{\min}(H_{t})\geqslant-\dfrac{1}{2}\sqrt{L_{2}\epsilon}$, Lemma 2.4 and (3.4), we have
$\|\nabla P(x_{t})\|\leqslant\|\nabla P(x_{t})-g_{t}\|+\|g_{t}\|\leqslant\frac{37}{36}\epsilon,$
and
$\nabla^{2}P(x_{t})\succeq H_{t}-\|\nabla^{2}P(x_{t})-H_{t}\|I\succeq\lambda_{\min}(H_{t})-\epsilon_{2}I\succeq-\frac{5}{9}\sqrt{L_{2}\epsilon}.$
The proof is then completed.
We are now ready to establish the iteration complexity for the LMNegCur algorithm. In particular, let $\epsilon>0$ be any given target accuracy. We denote the first iteration index to achieve an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point the same as (2.24). We further denote the first iteration index of the LMNegCur algorithm to achieve $\|g_{t}\|\leqslant\epsilon$ and $\lambda_{\min}(H_{t})\geqslant-\frac{1}{2}\sqrt{L_{2}\epsilon}$ by
$\bar{T}(\epsilon):=\min\{t|\|g_{t}\|\leqslant\epsilon~{}\text{and}~{}\lambda_{\min}(H_{t})\geqslant-\frac{1}{2}\sqrt{L_{2}\epsilon}\}.$
Then, from Lemma 3.3, we know that $x_{\bar{T}(\epsilon)}$ is already an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point, indicating that $T(\epsilon)\leqslant\bar{T}(\epsilon)$.
Let us define the following index sets to facilitate the complexity analysis,
$\mathcal{F}_{\bar{T}(\epsilon)}=\big{\{}t<\bar{T}(\epsilon):P(x_{t+1})-P(x_{t})\leqslant-\frac{1}{36\sqrt{L_{2}}}\max\{\|g_{t}\|^{3/2},\epsilon^{3/2}\}\big{\}},$ (3.17)
$\mathcal{G}_{\bar{T}(\epsilon)}=\left\{t<\bar{T}(\epsilon):\|g_{t+1}\|\leqslant\frac{1}{2}\|g_{t}\|,t\notin\mathcal{F}_{\bar{T}(\epsilon)}\right\}.$ (3.18)
From Lemmas 3.1 and 3.2, we conclude that each iteration $t$ belongs to one of the above two sets $\mathcal{F}_{\bar{T}(\epsilon)}$ and $\mathcal{G}_{\bar{T}(\epsilon)}$.
Similar to the proof in Section 2, we give an upper bound on $T(\epsilon)$ in the following theorem.
Theorem 3.1
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by Algorithm 2. Set
$\epsilon_{1}=\min\{\frac{1}{36},\frac{\sqrt{L_{2}}}{12L_{1}}\}\epsilon^{3/2},~{}\epsilon_{2}=\frac{\sqrt{L_{2}}}{18}\epsilon^{1/2},$ (3.19)
and choose $N_{t}$ in (2.8). We have
$T(\epsilon)\leqslant\lceil\log(\tilde{G}/\epsilon)/\log(2)\rceil|\mathcal{F}_{\bar{T}(\epsilon)}|+|\mathcal{F}_{\bar{T}(\epsilon)}|+1,$ (3.20)
where
$|\mathcal{F}_{\bar{T}(\epsilon)}|\leqslant 36\sqrt{L_{2}}(P(x_{0})-P^{*})\epsilon^{-3/2},$
$\tilde{G}:=\max\big{\{}\frac{7}{3}\tilde{C},1+\frac{L_{1}}{\sqrt{L_{2}}}\tilde{C}^{1/2}+\tilde{C},2+\frac{L_{1}}{\sqrt{L_{2}}}\big{\}}$ with $\tilde{C}:=\left(36\sqrt{L_{2}}\left(P(x_{0})-P^{*}\right)\right)^{2/3}=\mathcal{O}(\rho^{1/3}\kappa)$, which means that $\tilde{G}=\mathcal{O}(\rho^{1/3}\kappa+\ell\rho^{-1/3})$.
We omit the proofs here since they are almost identical to those in Theorem 2.1.
Remark 2
Theorem 3.1 shows that the iteration complexity of Algorithm 2 to find an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$ is $\tilde{\mathcal{O}}(\sqrt{L_{2}}\epsilon^{-1.5})=\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-1.5})$. Moreover, we can also have a rough estimate of the number of inner iterations, i.e., $N_{t}$. If $x_{t}$ is not an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point, according to (3.2) and (3.9), we have $\|s_{t}\|\geqslant\mathcal{O}(\min\{\sqrt{\frac{\epsilon}{L_{2}}},\frac{\epsilon}{L_{1}+\sqrt{L_{2}}}\})$. By further combining with Lemma 2.4, we obtain that $A=\mathcal{O}(\rho^{0.5}\kappa^{0.5}\ell^{-2}\epsilon^{1.5}+\rho^{-0.5}\kappa^{-0.5}\epsilon)$, $N_{1}=\mathcal{O}(\kappa\log\frac{\|y_{0}-y^{*}(x_{0})\|}{A})$ and $N_{t}=\mathcal{O}(\kappa\log\frac{A+\kappa\|s_{t-1}\|}{A})$ ($t\geqslant 2)$. Following the same discussion as in Remark 1, through the Nesterov’s accelerated gradient algorithm, the convergence speed can be further improved to $N_{1}=\mathcal{O}(\sqrt{\kappa}\log\frac{\sqrt{\kappa+1}\|y_{0}-y^{*}(x_{0})\|}{A})$ and $N_{t}=\mathcal{O}(\sqrt{\kappa}\log(\sqrt{\kappa+1}+\frac{\sqrt{\kappa+1}\kappa\|s_{t-1}\|}{A})$ ($t\geqslant 2$).
4 Numerical Results
In this section, we compare the numerical performance of the proposed GRTR and LMNegCur algorithms with the gradient descent (GDA) algorithm, the MCN algorithm Luo2022FindingSS , the MINIMAX-TR algorithm, and the MINIMAX-TRACE algorithm Yao2024TwoTR in solving a synthetic minimax optimization problem and an adversarial deep learning problem. The first experiment is implemented using Python 3.6 on a laptop with an Intel Core i7 2.8GHz and 8GB of RAM, while the second experiment is conducted on a 2.2GHz CPU, 80GB of RAM, and an Nvidia RTX 4090D GPU.
4.1 A synthetic minimax optimization problem
We consider the following nonconvex-strongly concave minimax problem Yao2024TwoTR :
$\min\limits_{x\in\mathbb{R}^{n}}\max\limits_{y\in\mathbb{R}^{m}}g(x)-\frac{1}{2}y^{2},$ (4.1)
where
$g(x)=\begin{cases}g_{i,1}(x)&x_{1},\cdots,x_{i-1}\in[2\tau,6\tau],x_{i}\in[0,\tau],x_{i+1},\cdots,x_{n}\in[0,\tau],\\ &1\leqslant i\leqslant n-1;\\ g_{i,2}(x)&x_{1},\cdots,x_{i-1}\in[2\tau,6\tau],x_{i}\in[\tau,2\tau],x_{i+1},\cdots,x_{n}\in[0,\tau],\\ &1\leqslant i\leqslant n-1;\\ g_{n,1}(x)&x_{1},\cdots,x_{n-1}\in[2\tau,6\tau],x_{n}\in[0,\tau];\\ g_{n,2}(x)&x_{1},\cdots,x_{n-1}\in[2\tau,6\tau],x_{n}\in[\tau,2\tau];\\ g_{n+1,1}(x)&x_{1},\cdots,x_{n}\in[2\tau,6\tau],\end{cases}$
with
$g_{i,1}(x)=\sum_{j=1}^{i-1}L(x_{j}-4\tau)^{2}-\gamma x_{i}^{2}+\sum_{j=i+1}^{n}Lx_{j}^{2}-(i-1)\nu,1\leqslant i\leqslant n-1,$
$g_{i,2}(x)=\sum_{j=1}^{i-1}L(x_{j}-4\tau)^{2}+h(x)+\sum_{j=i+2}^{n}Lx_{j}^{2}-(i-1)\nu,1\leqslant i\leqslant n-1,$
$g_{n,1}(x)=\sum_{j=1}^{n-1}L(x_{j}-4\tau)^{2}-\gamma x_{n}^{2}-(n-1)\nu,$
$g_{n,2}(x)=\sum_{j=1}^{n-1}L(x_{j}-4\tau)^{2}+h(x)-(n-1)\nu,$
$g_{n+1,1}(x)=\sum_{j=1}^{n}L(x_{j}-4\tau)^{2}-n\nu,$
and
$h(x)=\begin{cases}h_{1}(x_{i})+h_{2}(x_{i})x_{i+1}^{2}&x_{1},\cdots,x_{i-1}\in[2\tau,6\tau],\ x_{i}\in[\tau,2\tau],\\ &x_{i+1},\cdots,x_{n}\in[0,\tau],1\leqslant i\leqslant n-1;\\ h_{1}(x_{n})&x_{1},\cdots,x_{n-1}\in[2\tau,6\tau],\ x_{n}\in[\tau,2\tau];\\ 0&\text{otherwise},\end{cases}$
where
$h_{1}(x)$ $=-\gamma x^{2}+\frac{(-14L+10\gamma)(x-\tau)^{3}}{3\tau}+\frac{(5L-3\gamma)(x-\tau)^{4}}{2\tau^{2}},$
$h_{2}(x)$ $=-\gamma-\frac{10(L+\gamma)(x-2\tau)^{3}}{\tau^{3}}-\frac{15(L+\gamma)(x-2\tau)^{4}}{\tau^{4}}-\frac{6(L+\gamma)(x-2\tau)^{5}}{\tau^{5}},$
and
$L>0,\ \gamma>0,\ \tau=e,\ \nu=-h_{1}(2\tau)+4L\tau^{2}.$
Note that $g(x)$ is only defined on the following domain:
$D_{0}=\bigcup_{i=1}^{n+1}\left\{x\in\mathbb{R}^{n}:6\tau\geqslant x_{1},\cdots,x_{i-1}\geqslant 2\tau,2\tau\geqslant x_{i}\geqslant 0,\tau\geqslant x_{i+1},\cdots,x_{n}\geqslant 0\right\}.$
By Lemma A.3 in Du2017GradientDC , we know that there is only one local optimum, i.e., $(4\tau,...,4\tau)^{\top}$ and $d$ stationary points, i.e.,
$(0,\cdots,0)^{\top},(4\tau,0,\cdots,0)^{\top},\cdots,(4\tau,\cdots,4\tau,0)^{\top}.$
In the experiment, we set the number of inner iterations $N_{t}$ to $1000$, the dimension of the variable $y$ to $5$, and the dimension of the variable $x$ to $\{10,20\}$. The initial point $x_{0}$ is set to $(10^{-3},\cdots,10^{-3})^{\top}$, close to the stationary point $(0,\cdots,0)^{\top}$, and $y_{0}$ is a random vector. For the problem (4.1), the parameter $L$ is $\{1,1.5,2\}$, and the parameter $\gamma$ is $1$. Figure 1 shows the numerical performance of the six tested algorithms for solving the problem (4.1) under different parameter selections, where the horizontal axis represents the running time and the vertical axis represents the difference between $P(x)$ and the optimal value $P^{*}$.
Since there are $n$ saddle points in problem (4.1), and it is difficult for the GDA algorithm to escape from the saddle point when it is trapped in the saddle point, this also causes the curve of the GDA algorithm to present multiple horizontal line segments. As can be seen from Fig. 1, all second-order algorithms can effectively escape from the saddle point. The MINIMAX-TR algorithm converges slowly because the trust region radius is set to a small value. The proposed GRTR algorithm and LMNegCur algorithm are both better than the state-of-the-art trust region algorithm MINIMAX-TRACE and the cubic regularization algorithm MCN. Numerical results show the effectiveness of the proposed GRTR algorithm and LMNegCur algorithm.
Figure 1: Numerical results of the six tested algorithms for solving (4.1) with different choices of $n$, $L$ and $\gamma$.
4.2 An adversarial deep learning problem
The adversarial deep learning problem aims to train a convolutional neural network model with strong adversarially-robust capabilities, which can be formulated as the following minimax optimization problem Chen2023ACR :
$\min_{x}\max_{y=\{y_{i}\}_{i=1}^{n}}\frac{1}{n}\sum_{i=1}^{n}\left[\ell(h_{x}(y_{i}),b_{i})-\lambda\|y_{i}-a_{i}\|^{2}\right],$ (4.2)
where $x$ is the parameter of the neural network $h_{x}(\cdot)$, $(a_{i},b_{i})$ corresponds to the $i$th image and label respectively, and $y_{i}$ represents the adversarial sample corresponding to $a_{i}$.
In the experimental setting, we choose the cross entropy loss function as $\ell$ and the penalty coefficient as $\lambda=2$. We use the MNIST dataset Yann1998GradL , which contains 50k training samples and 10k test samples. The network model we use is a convolutional neural network consisting of a convolutional block and a fully connected layer. Specifically, the convolutional block contains a convolutional layer and a sigmoid activation layer. The convolutional layer has $1$ input channel and $1$ output channel, a kernel size of $3$, a stride of $4$, and a padding of $1$. The input dimension of the connection layer is $49$ and the output dimension is $10$.
We compare the proposed GRTR and LMNegCur algorithms with the GDA, MINIMAX-TR, and MCN algorithms. For all algorithms, we set the batch size to $64$. We select the gradient Lipschitz constant $L_{1}$ from $\{1,5,10\}$ and the Hessian Lipschitz constant $L_{2}$ from $\{10,50,100\}$ for experiments. The experimental results showing the most stable and optimal performance are then shown. For the GRTR algorithm, we set $N_{t}=30$, $\eta_{y}=0.01$, $\sigma=\sqrt{10}$, $r=1/\sqrt{10}$, and $\epsilon=0.01$, and solve the trust region subproblem by iterating $10$ times Curtis2021TrustN the truncated conjugate gradient (CG) method. For the LMNegCur algorithm and the MINIMAX-TR algorithm, we set $N_{t}=30$, $\eta_{y}=0.01$, $L_{2}=50$, and $\epsilon=0.01$. For the GDA algorithm, we set the step size of both $x$ and $y$ to 0.01. In addition, for the MCN algorithm, we solve the subproblem about ${y}$ by $30$ gradient descent iterations with a step size of $0.01$, which is the same as the GRTR algorithm and the LMNegCur algorithm. In addition, the cubic subproblem is solved by Algorithm 4 in Luo2022FindingSS , where $L=2$, $M=50$, $\sigma=0.001$, and $\mathcal{K}(\epsilon,\delta^{{}^{\prime}})=10$.
Figure 2 shows the numerical performance of the five tested algorithms when solving problem (4.2), where in sub-figure (a), the horizontal axis represents the running time and the vertical axis represents the test accuracy; while in sub-figure (b), the horizontal axis represents the number of iterations $t$ and the vertical axis represents the function value $P(x)$. It can be seen that the number of iterations and convergence time of all second-order algorithms are significantly less than that of the GDA algorithm, while GRTR performs slightly better than other second-order algorithms. These results show the advantages of the two proposed algorithm.
Figure 2: Numerical results of the five tested algorithms for solving (4.2).
5 Conclusions
In this paper, we propose a gradient norm regularized trust region (GRTR) algorithm and a Levenberg-Marquardt algorithm with a gradient norm regularization coefficient and use the negative curvature direction to correct the iteration direction (LMNegCur). Both the GRTR algorithm and the LMNegCur algorithm can achieve an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point with an iteration complexity of $\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-1.5})$. Extending the proposed algorithms to achieve optimal iteration complexity but with inexact subproblem solutions deserves further investigation.